# Intrinsic training dynamics of deep neural networks

### v1

@@ Conclusion
Since $t$ is fixed, in the following we lighten notations by dropping it from the equations. The proof of (47) is direct by showing that $\Lambda_{s}$ and $(Z_{s}^{-1})^{\top}Z_{1}^{\top}\nabla f(Z_{1})$ satisfy the same ODE (39) with the same value at $s=1$. Thus we only need to show (46).
@@
To show (46) we will use Riesz theorem to identify the expression of the gradient. We thus will consider the Hilbert space
@@
We recall that $Z_{s}(\theta)$ is the unique solution of the state equation (17):
@@
where $B_{u}^{\epsilon}\coloneqq\frac{\mathcal{A}_{u}^{\varepsilon}-\mathcal{A}_{u}}{\varepsilon}=\frac{\mathcal{A}_{u}^{\varepsilon}-\mathcal{A}_{u}^{0}}{\varepsilon}$ satisfies $B_{u}^{\varepsilon}\underset{\varepsilon\rightarrow 0}{\longrightarrow}\left.\frac{d}{d\varepsilon}\mathcal{A}_{u}^{\varepsilon}\right|_{\varepsilon=0}=\delta\mathcal{A}_{u}$ (as $\epsilon\mapsto\mathcal{A}_{u}^{\varepsilon}$ is $\mathcal{C}^{1}$) and where $\varepsilon\in[0,1]\mapsto B_{u}^{\varepsilon}$ is continuous (at 0, we define $B_{u}^{0}=\delta\mathcal{A}_{u}$) as $\varepsilon\mapsto\mathcal{A}_{u}^{\varepsilon}$ is $\mathcal{C}^{1}$), and thus is bounded on $[0,1]$ by a constant that does not depend on $\varepsilon$. By dominated convergence, when $\varepsilon\rightarrow 0$ in (51) one obtains the limit:
@@
which coincides with the unique solution of (50).
@@
Using (53) with $M=Z_{1}(\theta)$ and $H=\delta Z_{1}$,
@@
By inserting the expression of $\delta Z_{1}$ from (52) into (54), one has:
@@
The transpose in (55) finally yields the required formula. ∎
@@
In particular if $\theta_{L}$ satisfies the relaxed balanced condition
@@
Appendix M Proof of Theorem˜4.8.
@@
By (16) and the relation (46) (shown in Section˜L.2) we have for any $s\in[0,1]$
@@
Similarly as $U_{j-1}\cdots U_{0}=X_{j}=Z_{s_{j}}+\mathcal{O}(h)$ by (56), we get $Z_{s_{j}}^{\top}=(U_{j-1}\cdots U_{0})^{\top}+\mathcal{O}(h)$ hence
@@
As $\lambda(\cdot)$ is continuous, $C_{\lambda}\leq\|\lambda\|_{\infty}<\infty$ for any $L$. Similarly, we already used that as $s\in[0,1]\mapsto\mathcal{A}_{s}$ is continuous, $C_{U}=1+\mathcal{O}(h)$, and thus $C_{U}^{2}=1+\mathcal{O}(h)$, again with implicit constant independent of $L$. Moreover by (M.1), $\eta h^{2}=\eta/L^{2}=o(h^{2})$, and we obtain $C_{1}=(C_{U}^{2}+\eta h^{2}-1)/h=(\mathcal{O}(h)+o(h^{2}))/h=\mathcal{O}(1)$, hence $C_{1}$ is bounded uniformly. Finally we obtain
@@
and use the shorthand $A_{k}:=A_{k}^{L}(t)\coloneqq\mathcal{A}_{s_{k}}(t)\in\mathbb{R}^{n\times n}$, for $0\leq k\leq L-1$. Since $U_{k}=\mathrm{Id}_{n}+hA_{k}$ with $h=1/L$ and $s_{L-1}-s_{i}=\frac{L-1}{L}-\frac{i}{L}=1-\frac{i+1}{L}$ for each integer $i$, we have (using Riemann integration as $\lambda(\cdot)$ is continuous): since all the matrices in the product (66) commute
@@
We denote $\psi_{1}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(1-v)\mathrm{d}v\mathrm{d}u$. By (61) and the above derivations one has
@@
We denote $\psi_{2}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(v)\mathrm{d}v\mathrm{d}u$. By (68) and the above derivations we have
@@
Finally, combining (61)-(65)-(66)-(67) and (68)-(69)-(70), we obtain
@@
$\partial_{t}Z_{1}(t)={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-}\int_{0}^{1}(Z_{1}Z_{1}^{\top})^{1-s}\exp(\gamma(s))\nabla f(Z_{1})(Z_{1}^{\top}Z_{1})^{s}\mathrm{d}s,$
@@
M.2 Proof of (56)-(57)
We now show that  (56)-(57) hold for any $t$.
@@
Because the right-hand side of (72) $(s,W)\mapsto\begin{pmatrix}\mathcal{A}_{s}(t)&0\\[4.0pt] \partial_{t}\mathcal{A}_{s}(t)&\mathcal{A}_{s}(t)\end{pmatrix}\!W$ is $\mathcal{C}^{1}$ (indeed both $s\mapsto\mathcal{A}_{s}(t)$ and $s\mapsto\partial_{t}\mathcal{A}_{s}(t)$ are $C^{1}$ (cf Theorem˜L.1) for each $t$), the Euler explicit scheme converges at order one (see e.g. (Berthelin, 2017, Proposition 10.30)):

### v2

@@ F.2.2 Rank properties
Since $t$ is fixed, in the following we lighten notations by dropping it from the equations. The proof of equation 49 is direct by showing that $\Lambda_{s}$ and $(Z_{s}^{-1})^{\top}Z_{1}^{\top}\nabla f(Z_{1})$ satisfy the same ODE equation 41 with the same value at $s=1$. Thus we only need to show equation 48.
@@
To show equation 48 we will use Riesz theorem to identify the expression of the gradient. We thus will consider the Hilbert space
@@
We recall that $Z_{s}(\theta)$ is the unique solution of the state equation equation 16:
@@
where $B_{u}^{\epsilon}\coloneqq\frac{\mathcal{A}_{u}^{\varepsilon}-\mathcal{A}_{u}}{\varepsilon}=\frac{\mathcal{A}_{u}^{\varepsilon}-\mathcal{A}_{u}^{0}}{\varepsilon}$ satisfies $B_{u}^{\varepsilon}\underset{\varepsilon\rightarrow 0}{\longrightarrow}\left.\frac{d}{d\varepsilon}\mathcal{A}_{u}^{\varepsilon}\right|_{\varepsilon=0}=\delta\mathcal{A}_{u}$ (as $\epsilon\mapsto\mathcal{A}_{u}^{\varepsilon}$ is $\mathcal{C}^{1}$) and where $\varepsilon\in[0,1]\mapsto B_{u}^{\varepsilon}$ is continuous (at 0, we define $B_{u}^{0}=\delta\mathcal{A}_{u}$) as $\varepsilon\mapsto\mathcal{A}_{u}^{\varepsilon}$ is $\mathcal{C}^{1}$), and thus is bounded on $[0,1]$ by a constant that does not depend on $\varepsilon$. By dominated convergence, when $\varepsilon\rightarrow 0$ in equation 53 one obtains the limit:
@@
which coincides with the unique solution of equation 52.
@@
Using equation 55 with $M=Z_{1}(\theta)$ and $H=\delta Z_{1}$,
@@
By inserting the expression of $\delta Z_{1}$ from equation 54 into equation 56, one has:
@@
The transpose in equation 57 finally yields the required formula. ∎
@@
In particular if $\theta_{L}$ satisfies the quasi balanced condition
@@
Appendix M Proof of Theorem˜3.13.
@@
By equation 15 and the relation equation 48 (shown in Section˜L.2) we have for any $s\in[0,1]$
@@
Similarly as $U_{j-1}\cdots U_{0}=X_{j}=Z_{s_{j}}+\mathcal{O}(h)$ by equation 58, we get $Z_{s_{j}}^{\top}=(U_{j-1}\cdots U_{0})^{\top}+\mathcal{O}(h)$ hence
@@
As $\lambda(\cdot)$ is continuous, $C_{\lambda}\leq\|\lambda\|_{\infty}<\infty$ for any $L$. Similarly, we already used that as $s\in[0,1]\mapsto\mathcal{A}_{s}$ is continuous, $C_{U}=1+\mathcal{O}(h)$, and thus $C_{U}^{2}=1+\mathcal{O}(h)$, again with implicit constant independent of $L$. Moreover by equation M.1, $\eta h^{2}=\eta/L^{2}=o(h^{2})$, and we obtain $C_{1}=(C_{U}^{2}+\eta h^{2}-1)/h=(\mathcal{O}(h)+o(h^{2}))/h=\mathcal{O}(1)$, hence $C_{1}$ is bounded uniformly. Finally we obtain
@@
and use the shorthand $A_{k}:=A_{k}^{L}(t)\coloneqq\mathcal{A}_{s_{k}}(t)\in\mathbb{R}^{n\times n}$, for $0\leq k\leq L-1$. Since $U_{k}=\mathrm{Id}_{n}+hA_{k}$ with $h=1/L$ and $s_{L-1}-s_{i}=\frac{L-1}{L}-\frac{i}{L}=1-\frac{i+1}{L}$ for each integer $i$, we have (using Riemann integration as $\lambda(\cdot)$ is continuous): since all the matrices in the product equation 68 commute
@@
We denote $\psi_{1}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(1-v)\mathrm{d}v\mathrm{d}u$. By equation 63 and the above derivations one has
@@
We denote $\psi_{2}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(v)\mathrm{d}v\mathrm{d}u$. By equation 70 and the above derivations we have
@@
Finally, combining equation 63-equation 67-equation 68-equation 69 and equation 70-equation 71-equation 72, we obtain
@@
$\partial_{t}Z_{1}(t)=-\int_{0}^{1}(Z_{1}Z_{1}^{\top})^{1-s}\exp(\gamma(s))\nabla f(Z_{1})(Z_{1}^{\top}Z_{1})^{s}\mathrm{d}s,$
@@
M.2 Proof of equation 58-equation 59
We now show that  equation 58-equation 59 hold for any $t$.
@@
Because the right-hand side of equation 74 $(s,W)\mapsto\begin{pmatrix}\mathcal{A}_{s}(t)&0\\[4.0pt] \partial_{t}\mathcal{A}_{s}(t)&\mathcal{A}_{s}(t)\end{pmatrix}\!W$ is $\mathcal{C}^{1}$ (indeed both $s\mapsto\mathcal{A}_{s}(t)$ and $s\mapsto\partial_{t}\mathcal{A}_{s}(t)$ are $C^{1}$ (cf Theorem˜L.1) for each $t$), the Euler explicit scheme converges at order one (see e.g. (Berthelin, 2017, Proposition 10.30)):
@@
Appendix N LLM Usage
The authors of this paper used Large Language Models to aid and polish the writing of this paper and as a tool to make some of the proofs.